The half-life of strontium-91, 91

38Sr is 9.70 h. Find 

(a) its decay constant, and

(b) for an initial 1.00-g sample, the activity after 15 hours. 


1
Expert's answer
2021-11-02T10:01:45-0400

a)

"\lambda=\frac{\ln 2}{T_{1/2}}=19.8\cdot 10^{-6}~s^{-1},"

b)

"A=\lambda \frac mM N_A 2^{-\frac t{T_{1/2}}}=44.8\cdot 10^{15}~Bq."